OpenStudy (anonymous):

[(y^2+y) / (3xy)] * [(3x+3) / (y^2+2y+1)] Goodness; guide me?

OpenStudy (anonymous):

\[\frac{y^2+y}{3xy} \cdot \frac{3x + 3}{y^2 + 2y + 1}\]\[=\frac{(y^2+y)(3x+3)}{3xy(y^2 + 2y + 1)}\] I'd also look to see what you can factor out of those factors in the numerator cause some of it will cancel.

OpenStudy (anonymous):

Multiply and Expand ((3 x+3) (y^2+y))/(3 x y (y^2+2 y+1)) Now, Simplify (x+1)/(x y+x) (x+1)/(x(y+1))

OpenStudy (anonymous):

Also that denominator can factor nicely

OpenStudy (anonymous):

polpak; do you just distribute all that? Thank youu.

OpenStudy (anonymous):

Don't distribute yet! See what you can factor out and cancel (for example a 3 factors out of the (3x+3) and cancels with the 3 on the bottom. And a y factors out of the \((y^2 +y)\) and will cancel too

OpenStudy (anonymous):

Only distribute after you are sure that nothing else can factor and cancel.

OpenStudy (anonymous):

Alright, thanks!